OpenStudy (anonymous):

factor: 26x^2 - 159x - 65

OpenStudy (campbell_st):

(2x - 13)(13x + 5)

OpenStudy (anonymous):

wow, that was fast... how did you get that??

OpenStudy (anonymous):

find factors of -159??

OpenStudy (campbell_st):

I looked at 26 and 65... and can see 13 as a common factor so assumed 13 x 2 and 13 x 5 and 13^2 = 169... just a case of putting it together

OpenStudy (campbell_st):

-159 = -13^2 +10

OpenStudy (anonymous):

wow thank you.. it just seems so backwards to me.. I can work the quad. function no problem but factoring.. just confuses me to no end

OpenStudy (campbell_st):

well initially I thought about the general quadratic formula, which will give the factors... then saw 13 and 5 with 2 and 13

OpenStudy (anonymous):

my teacher said to find the factors of a number that will add up to the sum of two numbers in the equation... I just get confused as to which number I need to find prime factors to

OpenStudy (anonymous):

Would this be the correct way in beginning to factor out the problem? 26x2 - 159x – 65 2x2 - 132 + 10 – 5

OpenStudy (anonymous):

Correction: 26x^2 - 159x - 65 2x^2 - 13^2 + 10 -5

OpenStudy (anonymous):

and then? 2x^2 - 13^2 + 5

OpenStudy (anonymous):

I worked out (2x - 13)(13x + 5) as FOIL.. and it's not adding up right...

OpenStudy (campbell_st):

2x * 13x = 26x^2 -13 *13x = -169x &&& like terms and add to - 159x 2x * 5 = 10x &&& -13*5 = 65

OpenStudy (anonymous):

(2x – 13)(13x+5) 26x2 + 10x – 159x – 65

OpenStudy (anonymous):

(13x + 5) (2x – 13) 26x2 – 169x + 10x – 65 26x2 – 159x - 65 FIRST, OUTSIDE, INSIDE, LAST

OpenStudy (anonymous):

just wish I knew how Luis got to (13x + 5) (2x – 13)